If a complement is 24 more than twice the angle what is the complement
Answer by Theo(13342)

x is your angle.

90 - x is the complement of your angle.

the formula for the complement of the angle being 24 more than twice the angle is shown below:

90 - x = 2 * x + 24

add x to both sides of the equation and subtract 24 from both sides of the equation and you get:

90 - 24 = 3 * x

simplify to get:

66 = 3 * x

solve for x to get x = 22.

your angle is 22.
the complement is 90 - 22 = 68.

the complement is 24 more than twice the angle becomes:
68 = 2 * 22 + 24 which becomes:
68 = 44 + 24 which becomes:
68 = 68 which is true.

your solution is that the complement of the angle is 68.